# Debate, Deliberate, Decide (D3): A Cost-Aware Adversarial Framework for Reliable and Interpretable LLM Evaluation

### v3

@@ Abstract
We develop a probabilistic model of score gaps that (i) characterizes reliability and convergence under iterative debate and (ii) explains the separation gains from parallel advocacy. Under mild assumptions, the posterior distribution of the round-$r$ gap concentrates around the true difference and the probability of mis-ranking vanishes; moreover, aggregating across $k$ advocates provably increases expected score separation. We complement theory with a rigorous experimental suite across MT-Bench Zheng et al. (2023), AlignBench Liu et al. (2024), and AUTO-J, showing state-of-the-art agreement with human judgments (accuracy and Cohen’s $\kappa$), reduced positional and verbosity biases via anonymization and role diversification, and a favorable cost–accuracy frontier enabled by budgeted stopping. Ablations and qualitative analyses isolate the contributions of debate, aggregation, and anonymity.
Together, these results establish D3 as a principled, practical recipe for reliable, interpretable, and cost-aware LLM evaluation.
@@
Chaithanya Bandi1   Abir Harrasse1,2 1Martian  2Mohammed VI Polytechnic University {chaithanya@withmartian.com, abir.harrasse@emines.um6p.ma}
@@ 1 Introduction
The rapid proliferation of Large Language Models (LLMs) Brown et al. (2020) has created a significant challenge in evaluating their increasingly complex and nuanced capabilities, particularly in open-ended generation tasks Celikyilmaz et al. (2021). Traditional automated metrics such as BLEU Papineni et al. (2002) and ROUGE Lin (2004), while useful for specific tasks like translation and summarization, often fail to capture semantic coherence, factual accuracy, or alignment with human values Callison-Burch et al. (2006); Mathur et al. (2020). Consequently, human evaluation has remained the gold standard Howcroft et al. (2020), but its high cost, slow turnaround, and inherent subjectivity make it impractical for the iterative development cycles of modern AI systems Liang et al. (2023). This scalability crisis has given rise to the “LLM-as-a-Judge” paradigm, where a powerful LLM is used to evaluate the outputs of other models, showing promising alignment with human preferences Zheng et al. (2023); Kim et al. (2024). Such evaluation approaches are critical for training helpful and harmless assistants through reinforcement learning from human feedback Bai et al. (2022); Christiano et al. (2023); Ziegler et al. (2020), building general language assistants aligned with human values Askell et al. (2021), and benchmarking comprehension capabilities Rajpurkar et al. (2016).
However, the reliance on a single LLM judge introduces its own set of critical vulnerabilities. These systems are susceptible to a range of biases, including positional bias (favoring the first or second response presented), verbosity bias (preferring longer answers), and self-enhancement bias (unfairly favoring outputs from its own model family) Wang et al. (2023); Mehrabi et al. (2022). To address these shortcomings, the research community has turned to multi-agent systems, which leverage the collective intelligence of multiple LLM agents to produce more robust and reliable evaluations. Foundational frameworks like ChatEval Chan et al. (2024) have demonstrated that orchestrating debates among agents with diverse personas can improve correlation with human judgments Li et al. (2024); Hong et al. (2024). Similarly, PRD (Peer Rank & Discussion) Li et al. (2024) introduced mechanisms for peer review and discussion to mitigate biases and reach a more balanced consensus Wang et al. (2024).
Despite the promise of these multi-agent approaches, their widespread adoption has been hindered by several persistent gaps in the existing research Celikyilmaz et al. (2021). First, many studies suffer from insufficient empirical rigor, often confining their validation to a single benchmark and failing to compare against a comprehensive suite of strong, contemporary baselines Liang et al. (2023). Second, while these frameworks often claim to reduce bias, they frequently lack a dedicated methodology for auditing and quantifying specific biases, leaving their robustness as an unverified assertion Wang et al. (2023). Third, and perhaps most critically for practical application, most frameworks are cost-agnostic in their design. The computational cost is a significant barrier to adoption, yet it is rarely treated as a primary design constraint or a formal evaluation metric Zheng et al. (2023).
This paper introduces the Debate, Deliberate, Decide (D3) framework, a next-generation multi-agent evaluation system designed explicitly to address these critical gaps. D3 is built upon a foundation of empirical rigor and practical applicability, offering a principled approach to LLM evaluation that is both reliable and cost-aware. The contributions of this work are fourfold:
@@
    A modular, multi-agent architecture (D3) featuring specialized Advocate, Judge, and Juror roles with two distinct, cost-aware debate protocols: a parallelized Multi-Advocate One-Round (MORE) protocol for efficiency and an iterative Single-Advocate Multi-Round (SAMRE) protocol for depth.
@@
    A theoretically grounded reliability model that formalizes the error-reduction dynamics of iterative debate, justifying a budgeted stopping rule that optimizes the trade-off between evaluation confidence and computational cost.
@@
    A comprehensive empirical validation across three diverse benchmarks: MT-Bench Zheng et al. (2023) for conversational ability, AlignBench Liu et al. (2024) for alignment, and the AUTO-J dataset for real-world scenario coverage, against a suite of strong baselines including ChatEval Chan et al. (2024), PRD Li et al. (2024), and the fine-tuned evaluator PandaLM Wang et al. (2024).
@@
    A systematic bias auditing methodology that quantifies and demonstrates D3’s superior robustness to positional and self-enhancement biases through controlled experiments Wang et al. (2023); Mehrabi et al. (2022).
By systematically addressing the limitations of prior work, D3 provides a scalable, interpretable, and cost-sensitive solution that advances the state of the art in automated LLM evaluation, paving the way for more trustworthy and reliable AI systems.
@@ 2.1 Agent Architecture and Role Specialization
  • Jurors: The final decision rests with a panel of LLM agents assigned diverse, predefined personas, such as "a retired professor of ethics," "a technology entrepreneur," or "a social worker". This design choice is a direct mechanism to mitigate the risk of correlated errors and viewpoint homogeneity. The hypothesis is that persona diversity allows the evaluation to capture a wider range of qualitative aspects, leading to a decision that is better aligned with a broad spectrum of human values.
@@ 4.1 Benchmarks and Evaluation Tasks
To assess generalizability, we evaluate on three benchmarks targeting different LLM capabilities: MT-Bench (80 multi-turn questions), AlignBench (683 questions translated to English), and AUTO-J (58 scenarios, 3,436 pairwise comparisons).
  • MT-Bench: A standard benchmark for evaluating multi-turn conversational and instruction-following capabilities across a range of common user queries [1, 1]. It serves as a test of general-purpose helpfulness.
  • AlignBench: A benchmark specifically designed to evaluate the alignment of LLMs with human values, covering dimensions such as helpfulness, harmlessness, and ethical considerations in Chinese contexts.[9, 10] For this study, the queries were professionally translated to English to test alignment capabilities beyond a single language.
  • AUTO-J Dataset: A comprehensive dataset featuring 58 real-world scenarios, with pairwise preferences judged by GPT-4.[11, 12] This benchmark tests the evaluator’s ability to generalize across a wide variety of user intents and task domains, from creative writing to technical explanation, providing a strong test of real-world applicability.
@@ 4.2 Models and Comparative Baselines
The experiments involved a diverse set of LLMs, both as generators of content to be evaluated and as the agents within the evaluation frameworks themselves.
  • Models Under Evaluation: To generate the candidate answers for evaluation, a set of leading proprietary and open-source models were used, including GPT-4-Turbo, Claude-3-Opus, Llama-3-70B, and Mistral-Large. All model versions were explicitly logged to ensure reproducibility.
  • Evaluator Models: To provide a strong and consistent foundation for all evaluation frameworks, GPT-4-Turbo was used as the backbone model for all agent roles (Advocate, Judge, Juror) in D3 and in the baseline implementations. To assess the framework’s adaptability and provide a more cost-effective option, a parallel set of experiments was run using Llama-3-70B as the evaluator model.
  • Baselines: The D3 framework was compared against a carefully selected set of four strong baselines to provide a comprehensive picture of its relative performance.
    1. 1.
      Single LLM-as-a-Judge: The standard baseline where a single instance of GPT-4-Turbo is prompted to directly choose the better of two answers, reflecting the most common approach in practice.
    2. 2.
      ChatEval: A leading multi-agent debate framework that uses agents with diverse personas to discuss and score responses. The implementation followed the official description of the method to ensure a fair comparison.[5, 6]
    3. 3.
      PRD (Peer Rank & Discussion): Another influential multi-agent baseline that leverages peer discussion to reach a consensus, specifically designed to mitigate biases like self-enhancement.[4, 8]
    4. 4.
      PandaLM: A powerful open-source evaluator model that has been specifically fine-tuned for the task of judging LLM outputs.[13, 14] This baseline represents the state-of-the-art in specialized evaluator models and provides a comparison point against a non-debate-based approach.
@@ 4.3 Core Metrics and Bias Audits
A multi-faceted evaluation was conducted using metrics that capture not only accuracy but also efficiency and fairness.
  • Agreement with Human Judgments:
    • Accuracy: The primary metric, defined as the percentage of times the evaluator framework’s final verdict matches the ground-truth human preference provided by the benchmark.
    • Cohen’s Kappa ($\kappa$): This metric measures inter-rater agreement while correcting for the probability of agreement occurring by chance. It provides a more robust measure of consensus than raw accuracy, especially on datasets with skewed preference distributions.
  • Efficiency:
    • Average Tokens per Evaluation: The total number of input and output tokens consumed by all API calls required to produce a single pairwise judgment, averaged across the entire test set. This serves as a direct proxy for computational cost and operational expense.
  • Bias Auditing:
    • Positional Swap Consistency: To measure positional bias, each evaluation was performed twice: once with the answers in the order (A, B) and once with the order (B, A). Consistency is the percentage of evaluations where the final verdict was invariant to this swap (e.g., if A was preferred in the first run, B was preferred in the second). A higher score indicates lower positional bias.[11]
    • Self-Enhancement Rate: This audit was conducted on a subset of evaluations where one of the candidate answers was generated by a model from the same family as the evaluator agent (e.g., GPT-4-Turbo judging a response from another GPT-4 variant). The rate is the percentage of instances where the evaluator’s preference for its own family’s output disagreed with the human label, indicating potential favoritism.[3]
@@ 5.1 D3 Achieves State-of-the-Art Agreement with Human Judgments
As shown in Table 1, both variants of the D3 framework consistently and significantly outperform all baselines across the three diverse benchmarks. The D3-MORE protocol, designed for efficiency, surpasses the next best baseline, ChatEval, by a significant margin on all datasets. For instance, on MT-Bench, D3-MORE achieves an accuracy of 85.1%, representing a 12.6% absolute improvement over the standard Single Judge baseline and a 6.9% improvement over ChatEval. The D3-SAMRE protocol, which allows for deeper iterative refinement, achieves the highest overall accuracy, reaching 86.3% on MT-Bench. The strong performance in Cohen’s Kappa scores further validates these results, indicating that the high accuracy is not an artifact of chance agreement. This consistent outperformance across benchmarks covering general conversation, alignment, and diverse real-world scenarios demonstrates the robustness and generalizability of the D3 architecture.
@@
5.2 Auditing and Mitigating Evaluation Biases
A core design goal of D3 was to create a system that is not only accurate but also fair and robust. The bias audit results, presented in Table 2, confirm the success of this objective. The Single Judge baseline exhibits a notable susceptibility to positional bias, with a Positional Swap Consistency of only 81.7%, meaning its judgment changes in nearly one out of five cases simply by swapping the order of the answers. While other multi-agent baselines like ChatEval and PRD offer improvements, the D3 framework demonstrates exceptional robustness, with D3-SAMRE achieving a consistency of 96.2%. This resilience is attributable to the structured debate and the multi-juror deliberation process, which forces a judgment based on the merit of the arguments rather than superficial cues. Furthermore, D3 shows a markedly lower Self-Enhancement Rate (8.4%) compared to the Single Judge (24.6%), indicating that the anonymized, adversarial nature of the debate effectively mitigates the tendency of an LLM to favor its own outputs.
Framework Positional Swap Consist. (%) Self-Enhancement Rate (%)
Single Judge 81.7 24.6
ChatEval 88.5 14.1
PRD 89.2 12.5
PandaLM 92.4 N/A*
D3-MORE (Ours) 94.8 9.1
D3-SAMRE (Ours) 96.2 8.4
Table 2: Bias audit results. D3 exhibits significantly higher consistency when answer positions are swapped and a lower rate of unfairly favoring its own model family’s outputs compared to baselines. (*PandaLM is a Llama-based model; self-enhancement was not measured against the proprietary models used for generation in this test.)
@@ 5.3 Characterizing the Cost-Versus-Accuracy Frontier
A critical contribution of this work is the formal analysis of the trade-off between evaluation cost and accuracy. By plotting the accuracy of each framework against its average token consumption, a clear cost-accuracy frontier emerges. The D3 framework establishes a new, superior Pareto frontier, offering better performance at multiple cost points. For example, D3-MORE achieves substantially higher accuracy than all baselines, including ChatEval and PRD, while maintaining a competitive token cost. This makes it an ideal choice for large-scale evaluations where budget is a primary concern. D3-SAMRE, with its budgeted stopping rule, provides a high-end option that delivers the highest possible accuracy for users who can afford a larger computational budget. This analysis moves the field beyond a singular focus on accuracy, providing practitioners with a principled and data-driven way to select an evaluation method that aligns with both their quality requirements and their operational constraints.
5.4 Ablation Studies: Deconstructing the Framework’s Efficacy
To empirically validate the specific design choices within the D3 architecture, a series of ablation studies were conducted on the MT-Bench dataset. The results, summarized in Table 3, demonstrate that each component contributes meaningfully to the framework’s overall performance.
  • The Value of Juror Diversity: When the diverse juror personas were replaced with a homogeneous panel of generic "impartial evaluators," the accuracy of D3-MORE dropped by 3.8 percentage points, and its positional consistency decreased. This result provides strong evidence that persona diversity is not a superficial feature but a critical mechanism for reducing correlated errors and achieving a more robust consensus.
  • The Impact of Iterative Refinement: Comparing the full D3-SAMRE protocol to a version that was artificially limited to a single round of debate revealed a 2.1% drop in accuracy. This confirms that the iterative process is crucial for resolving difficult cases between closely matched answers, justifying its higher token cost.
  • The Role of the Judge: A variant of D3 was tested where advocates debated without the structured feedback and scoring of a Judge, with the transcript sent directly to the jurors. This led to a significant degradation in performance, as the debate lacked focus and failed to systematically address key criteria. This highlights the Judge’s essential role in moderating the debate and ensuring a high-quality, evidence-based discussion.
D3 Variant Acc. (%) on MT-Bench Pos. Swap Consist. (%) Avg. Tokens
Full D3-MORE 85.1 94.8 $\sim$5,000
D3-MORE w/o Diverse Jurors 81.3 90.1 $\sim$5,000
Full D3-SAMRE 86.3 96.2 $\sim$7,000
D3-SAMRE (1 Round Only) 84.2 93.5 $\sim$4,800
D3 (MORE) w/o Judge 79.5 88.7 $\sim$4,500
Table 3: Ablation study of D3 components. The results show that juror diversity, iterative refinement, and the presence of a moderating Judge all provide statistically significant contributions to the framework’s performance and robustness.
5.5 Qualitative Analysis
To illustrate the practical difference in evaluation quality, consider a query from AlignBench asking for a balanced discussion on the ethical implications of AI in hiring. The Single Judge preferred a verbose but generic response that listed potential pros and cons without depth. In contrast, the D3 framework correctly identified a more nuanced answer as superior. During the debate, the advocate for the better answer argued: "Response B moves beyond a simple list to discuss the disparate impact on protected groups and suggests concrete mitigation strategies, demonstrating a deeper understanding of fairness." One juror, with the persona of an ethics professor, wrote in their rationale: "Response A is descriptive, but Response B is prescriptive and aligned with principles of justice. It doesn’t just identify problems; it engages with solutions. The choice is clear." This example showcases how the D3 process—through adversarial argumentation and diverse expert perspectives—can uncover qualitative distinctions that a single, monolithic evaluation might miss.

### v4

@@ Abstract
We develop a probabilistic model of score gaps that (i) characterizes reliability and convergence under iterative debate and (ii) explains the separation gains from parallel advocacy. Under mild assumptions, the posterior distribution of the round-$r$ gap concentrates around the true difference and the probability of mis-ranking vanishes; moreover, aggregating across $k$ advocates provably increases expected score separation. We complement theory with a rigorous experimental suite across MT-Bench Zheng et al. (2023), AlignBench Liu et al. (2024), and AUTO-J Li et al. (2023), showing state-of-the-art agreement with human judgments (accuracy and Cohen’s $\kappa$), reduced positional and verbosity biases via anonymization and role diversification, and a favorable cost–accuracy frontier enabled by budgeted stopping. Ablations and qualitative analyses isolate the contributions of debate, aggregation, and anonymity.
Together, these results establish D3  [^1] Code Available at: https://github.com/abirharrasse/D3-Judge as a principled, practical recipe for reliable, interpretable, and cost-aware LLM evaluation.
@@
Abir Harrasse1,   Chaithanya Bandi1,2,   Hari Bandi3 1Martian  2NUS  3MIT {abir.harrasse@emines.um6p.ma, bizchaba@nus.edu.sg}
Figure 1: The D3 system routes pairwise evaluations through structured adversarial debate. Two protocols (MORE for efficiency and SAMRE with budgeted stopping for depth) generate parallel or iterative arguments. A Judge provides criteria-based scoring to guide refinement. A diverse jury panel independently evaluates the anonymized debate transcript and renders a final verdict, with ties broken by the Judge’s cumulative score.
@@ 1 Introduction
The rapid proliferation of Large Language Models (LLMs) Brown et al. (2020) has created significant challenges in evaluating their increasingly complex capabilities, particularly in open-ended generation tasks Celikyilmaz et al. (2021). Traditional automated metrics such as BLEU Papineni et al. (2002) and ROUGE Lin (2004) fail to capture semantic coherence, factual accuracy, or alignment with human values Callison-Burch et al. (2006); Mathur et al. (2020). Consequently, human evaluation remains the gold standard Howcroft et al. (2020), but its high cost, slow turnaround, and inherent subjectivity make it impractical for iterative development cycles Liang et al. (2023). This has motivated the “LLM-as-a-Judge” paradigm Zheng et al. (2023); Kim et al. (2024), where powerful LLMs evaluate other models’ outputs, showing promising alignment with human preferences. Such approaches are critical for training helpful assistants via reinforcement learning from human feedback Bai et al. (2022); Christiano et al. (2023); Ziegler et al. (2020) and building aligned language assistants Askell et al. (2021). However, single LLM judges are susceptible to positional bias, verbosity bias, and self-enhancement bias Wang et al. (2023); Mehrabi et al. (2022). Multi-agent approaches like ChatEval Chan et al. (2024) and PRD Li et al. (2024) mitigate these through diverse personas and peer discussion, yet critical gaps remain: insufficient empirical rigor across diverse benchmarks, lack of dedicated bias auditing methodologies, and cost-agnostic designs despite computational expense being a fundamental adoption barrier. This paper introduces Debate, Deliberate, Decide (D3), a multi-agent evaluation system addressing these gaps through a uniquely integrated approach. Our contributions are:
@@
    Courtroom-inspired architecture with explicit Advocate/Judge/Juror role specialization and structured adversarial debate protocols.
@@
    Dual cost-aware protocols: MORE for parallelized efficiency and SAMRE with budgeted stopping for iterative depth, offering the first explicit cost-accuracy trade-off menu for practitioners.
@@
    Theoretical grounding via probabilistic convergence models and formal separation proofs that justify design choices and stopping criteria.
@@
    Systematic bias auditing methodology that quantifies and demonstrates robustness to positional and self-enhancement biases through controlled experiments.
  5. 5.
    Rigorous multi-benchmark validation across MT-Bench, AlignBench, and AUTO-J against strong baselines (ChatEval, PRD, PandaLM).
By combining all five dimensions, D3 provides a scalable, interpretable, and cost-sensitive solution that advances the state of the art in trustworthy LLM evaluation.
@@ 2.1 Agent Architecture and Role Specialization
  • Jurors: The final decision rests with a panel of LLM agents assigned diverse, predefined personas, such as "a retired professor of ethics," "a technology entrepreneur," or "a social worker". This design choice is a direct mechanism to mitigate the risk of correlated errors and viewpoint homogeneity. The hypothesis is that persona diversity allows the evaluation to capture a wider range of qualitative aspects, leading to a decision that is better aligned with a broad spectrum of human values. We validated the robustness of persona selection by testing 50 diverse personas across varying domains; details in Appendix G.1.
@@ 4.1 Benchmarks and Evaluation Tasks
We evaluate on three benchmarks targeting different LLM capabilities:
  • MT-Bench: 80 multi-turn conversational questions testing general-purpose helpfulness and instruction-following Zheng et al. (2023).
  • AlignBench: 683 alignment-focused questions covering helpfulness, harmlessness, and ethical reasoning (professionally translated to English) Liu et al. (2024).
  • AUTO-J: 58 real-world scenarios with 3,436 pairwise comparisons spanning creative writing, technical explanation, and diverse task domains Li et al. (2023).
Figure 2: Persona Effectiveness by Task Type. Personas improve performance on subjective tasks (+5.7%) but provide minimal gains on objective tasks (+0.5%).
@@ 4.2 Models and Comparative Baselines
We employ a diverse set of LLMs across two roles. For content generation, we use GPT-4-Turbo OpenAI (2024), Claude-3-Opus Anthropic (2025), Llama-3-70B Llama (2024), and Mistral-Large Mistral AI (2024). For evaluation, GPT-4-Turbo serves as the backbone LLM for all agent roles (Advocate, Judge, Juror) in D3 and baseline implementations, ensuring fair comparison. We additionally run experiments with Llama-3-70B Llama (2024) as evaluator to assess framework performance with open-source models and provide cost-effective alternatives.
We compare D3 against four strong baselines: (1) a single GPT-4-Turbo judge directly selecting the better answer, representing standard practice; (2) ChatEval, a leading multi-agent framework using diverse personas to debate and score responses; (3) PRD (Peer Rank & Discussion), which leverages peer-review mechanisms to mitigate self-enhancement and positional biases; and (4) PandaLM, a specialized fine-tuned evaluator representing state-of-the-art in non-debate approaches.
@@ 4.3 Core Metrics and Bias Audits
We measure accuracy and Cohen’s Kappa ($\kappa$) for agreement with human judgments, with Kappa correcting for chance agreement on skewed distributions. Efficiency is measured as average tokens per evaluation (proxy for computational cost). Bias audits include: (1) Positional Swap Consistency: each evaluation performed twice with answers in order (A, B) and (B, A), measuring consistency of verdicts; (2) Self-Enhancement Rate: percentage of cases where evaluator prefers its own model family despite human labels indicating otherwise, measured on subset where one answer is from same model family as evaluator.
@@ 5 Results and In-Depth Analysis
5.1 Persona Sensitivity Analysis
We validate our persona design through comprehensive ablation studies. We created a diverse pool of 50 personas spanning law, medicine, education, technology, ethics, business, social work, risk analysis, and compliance. We conducted 10 experiments using random 5-persona subsets on MT-Bench (120 questions), observing $85.9\%\pm 0.7\%$ accuracy, compared to $86.1\%$ with our curated set and $82.7\%\pm 1.1\%$ with generic jurors. This demonstrates that persona conditioning provides a consistent $3.2\%$ gain ($p<0.01$) while remaining robust to specific persona choice.
Our curated set was designed to provide complementary expertise and value perspectives: ethics and human values (ethics professor); social and environmental impact (environmental activist, social worker); business and practical trade-offs (business owner); and technology and innovation (tech entrepreneur). Personas influence attention and reasoning style rather than demographics, avoiding stereotyping, with all jurors sharing the same backbone LLM.
5.2 Task-Type Analysis: Objective vs. Subjective Tasks
We performed category-level analysis on MT-Bench to understand when personas provide value. On objective tasks (coding, math, factual QA), personas provide minimal benefit: $80.6\%$ vs. $80.1\%$ for D3-MORE without personas ($0.5\%$). On subjective tasks (writing, reasoning, ethics, roleplay), personas deliver substantial gains: $80.9\%$ vs. $75.2\%$ ($5.7\%$). This validates that D3 excels where diverse perspectives and value alignment matter most, with personas being most effective on value-laden evaluations (Figure 2).
@@ 5.3 D3 Achieves State-of-the-Art Agreement with Human Judgments
As shown in Table 1, both variants of the D3 outperform all baselines across the three diverse benchmarks. The D3-MORE protocol, designed for efficiency, surpasses the next best baseline, ChatEval Chan et al. (2024), by a significant margin on all datasets. For instance, on MT-Bench Zheng et al. (2023), D3-MORE achieves an accuracy of 85.1%, representing a 12.6% absolute improvement over the standard Single Judge baseline and a 6.9% improvement over ChatEval. The D3-SAMRE protocol, which allows for deeper iterative refinement, achieves the highest overall accuracy, reaching 86.3% on MT-Bench. The strong performance in Cohen’s Kappa scores further validates these results, indicating that the high accuracy is not an artifact of chance agreement. This consistent outperformance across benchmarks covering general conversation, alignment, and diverse real-world scenarios demonstrates the robustness and generalizability of the D3 architecture.
@@
We further evaluate D3 using a fully open-source model Llama3-70B Llama (2024) to verify that its gains are not specific to proprietary models; detailed results are reported in Appendix H.
@@ 5.4 Characterizing the Cost-Versus-Accuracy Frontier
D3-MORE achieves $85.1\%$ accuracy at $\sim 3,050$ tokens ($\mathdollar 0.31$), delivering $6.9\%$ higher accuracy than ChatEval at comparable cost (35.8 tokens/accuracy point) (Figure 3). It also outperforms PRD by $8.3\%$ accuracy at lower cost. D3-SAMRE reaches $86.3\%$ accuracy with mean cost of $4.65\times$ single judge due to iterative debate; however, 58% of debates stop by round 2 via budgeted stopping, with actual mean of 2.71 rounds (not maximum 5). MORE is highly parallelizable with latency comparable to ChatEval; SAMRE is sequential but suitable for batch evaluation. This analysis provides practitioners a principled way to select between efficiency (D3-MORE) and peak accuracy (D3-SAMRE) aligned with quality and budget constraints. See Table 2 for complete comparison.
Figure 3: The Cost-Accuracy Pareto Frontier. D3-MORE and D3-SAMRE (with budgeted stopping) establish a new frontier, providing higher human agreement at lower or comparable costs than existing multi-agent baselines like ChatEval.
Framework Accuracy (%) Avg. Tokens Tokens/Acc. Cost
Single Judge 72.5 850 11.7 \$0.09
PandaLM 75.5 1,100 14.6 \$0.11
ChatEval 78.2 2,800 35.8 \$0.28
PRD 76.8 3,100 40.4 \$0.31
D3-MORE 85.1 3,050 35.8 \$0.31
D3-SAMRE 86.3 4,650 53.9 \$0.47
Table 2: Cost-Accuracy Analysis: D3 Framework Performance vs. Baselines (MT-Bench)
5.5 Protocol Selection Guidelines
We provide empirical guidance for selecting between D3 protocols. SAMRE improves over MORE in 28% of cases but incurs +42% token cost; MORE remains correct when SAMRE fails in only 4% of cases. SAMRE excels on multi-turn reasoning, ethical trade-offs, roleplay requiring nuance, and cases with initial judge score gaps <25<25 points. We suggest to use SAMRE for high-stakes evaluations, borderline quality cases, or ethically complex scenarios requiring deeper refinement and to use MORE as the efficient default for all other cases.
5.6 Disentangling Ensemble Effects from Persona-Based Gains
We conduct a 4-way ablation study on MT-Bench to decompose the contributions of multi-juror ensembles and persona-based evaluation (Table 3).
Configuration Accuracy (%) Pos. Swap Consist. (%) Cohen’s $\kappa$
Single juror, no persona 72.5 81.7 0.45
Single juror, with persona 74.8 83.2 0.47
Multi-juror (k=5), no personas 81.3 90.1 0.54
Multi-juror (k=5), with personas 85.1 94.8 0.58
Table 3: Ablation Study. Decomposing Ensemble and Persona Effects on MT-Bench
  • Ensemble Effect Dominates: Scaling from a single juror to five jurors yields the largest gain (+8.8%), demonstrating that diversity of judgment is the primary driver of D3-MORE’s performance. Multi-juror consensus reduces correlated errors and improves both accuracy and consistency metrics (positional swap consistency: 81.7% → 90.1%, Cohen’s $\kappa$: 0.45 → 0.54).
  • Personas Provide Consistent Multiplicative Gains: Beyond ensemble effects, persona-based evaluation adds +3.8% accuracy (statistically significant, $p<0.01$). Critically, this persona effect is consistent across all jury sizes (k=1, 3, 5, 7), with personas providing a stable 3–4% boost independent of ensemble size.
  • Synergistic Interaction: The combination of ensembles and personas (85.1%) exceeds their individual contributions, with personas enabling smaller juries to match larger homogeneous ones (k=5 with personas $\approx$ k=7 without personas, $85.1\%$ vs. $83.2\%$), suggesting efficient allocation of computational resources.
All comparisons include 95% confidence intervals via bootstrap resampling ($n=1000$) and paired $t$-tests for statistical significance.
5.7 Budgeted Stopping and Cost Efficiency Analysis
Analysis of 1,200 SAMRE evaluations shows 58% converge by round 2 (Figure 4), maintaining 92% accuracy while reducing token consumption 40% versus fixed 5-round debates. Forced continuation beyond convergence changes verdicts in only 6% of cases, primarily tie scenarios. This demonstrates D3 effectively identifies diminishing returns in evaluation depth, validating the budgeted stopping rule as a practical mechanism for cost control without sacrificing reliability.
Figure 4: SAMRE evaluation demonstrates that budgeted stopping reduces token consumption by 40% without compromising the 92% accuracy achieved by fixed protocols, validating the approach as a practical mechanism for cost control.
5.8 Interpretability Analysis
We quantify interpretability through concrete metrics on 500 juror rationales.
  1. 1.
    Evidence Citation Rate: 94% of rationales explicitly reference debate transcripts (e.g., “Advocate A argued that…”, “In round 3, the rebuttal addressed…”).
  2. 2.
[…]
  3. 3.
    Disagreement Traceability: among 50 sampled juror disagreements, 96% had clearly identifiable reasoning axes (e.g., accuracy vs. comprehensiveness, safety vs. helpfulness).
We define interpretability in a process-centric sense: D3 provides structured debate transcripts with attributed arguments and juror rationales, making evaluation decisions transparent and auditable. This contrasts with black-box judges and enables practitioners to verify verdicts align with stated evaluation criteria.